OpenStudy (anonymous):

3. Explain to Dr. Collier how to create a rough sketch of a graph of a fourth-degree polynomial function.

OpenStudy (kira_yamato):

So a fourth degree polynomial can be expressed as \[(x - \alpha)(x - \beta)(x - \gamma)(x - \delta) = 0, \alpha,\beta,\gamma,\delta \in \mathbb{C}\]

OpenStudy (anonymous):

o ok thanks those are the the x values and the y values are zero right

OpenStudy (kira_yamato):

So this will give you the roots to the graph of y = f(x). So you could easily graph them this way, where I'm assuming that the roots are real.